A station wagon leaves Seattle heading east at 40 mph. Three hours later, a Ferrari leaves Seattle on the same path as the station wagon, but going 95 mph. How many hours has the station wagon been traveling when it si caught by the ferrari?
Answer by checkley77(12844)

D=RT
D=40T FOR THE WAGON
D=95(T-3)
40T=95(T-3)
40T=95T-285
40T-95T=-285
-55T=-285
T=-285/-55
T=5.182 HOURS AFTER THE WAGON LEAVES THE FERRARI CATCHES UP.
40*5.182=95(5.182-3)
207.28=95*2.182
207.28~207.28